# Rethinking the handling of method failure in comparison studies

### v2

@@ 1 Introduction
In methodological research, comparison studies aim to compare the performance of methods to provide empirical evidence on their behavior and help data analysts choose appropriate methods for their setting. To be reliable, they require careful planning of the design, execution, and reporting of the results. These issues have gained increasing attention in recent years, both in the context of simulation studies based on artificially generated data (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024) and in the context of real data-based benchmarking studies (Boulesteix et al., , 2017; Friedrich and Friede, , 2024).
@@
In this work, we focus on a particular aspect of the design of comparison studies that affects both simulation and benchmark studies, namely “method failure.” Method failure occurs when a method under investigation fails to produce an output for a (real or simulated) data set considered in the comparison study. The issue of method failure in comparison studies has begun to attract attention in recent years (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024), often focusing on non-convergence. However, there is little guidance on how to handle method failure appropriately, whether it occurs as non-convergence or in another form. In particular, published reviews have shown that method failure goes unreported in the majority of cases. For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur(Hinds et al., , 2018). Even more notably, only $12$ of $85$ applicable research articles reviewed by Morris et al., (2019) report convergence as a performance measure. This contrasts with our perception—both through our own experiences and informal discussions with colleagues—that the problem of method failure is highly prevalent. In particular, signs of method failure are often traceable in the code provided for reproducibility, even when not reported in the manuscript. This is often clear through the use of explicit error-handling mechanisms (e.g., functions try() or tryCatch() in R or try-except constructs in Python).
@@
To illustrate the difficulty of handling method failure adequately, imagine a researcher conducting a comparison study in statistics or predictive modeling to compare the performance of three methods based on four simulation/benchmark data sets. In the context of classical statistics, they likely focus on comparing the methods’ performances (e.g., bias) in “estimating one or more population quantities” (Morris et al., , 2019), sometimes called “estimand(s),” and base the analysis on data simulated using a certain data-generating mechanism (DGM). In predictive modeling, they are likely to compare the methods’ predictive performances (e.g., accuracy), that is, their ability to make predictions on a new set of observations. Here, the analysis is typically performed on real benchmarking data sets. For simplicity, we refer to both settings as “simulation studies” and “benchmark studies,” respectively (though admitting that there may be intermediate situations, e.g., simulation studies investigating predictive performance on artificial data).
@@
Average accuracy ? 0.84 ?
Regardless of the form in which a researcher encounters method failure, it leads to NA for the associated method-data set combination where otherwise an estimate (for simulation studies) or a performance value (for benchmark studies) is expected. Table 1 depicts a corresponding summary of the results of the hypothetical researcher’s study across all simulation repetition/benchmark data sets and methods. In simulation studies, the NAs complicate the derivation of performance (e.g., bias as the mean deviation of the estimates from the true value across all repetitions). In benchmark studies, on the other hand, they hinder the aggregation of performance values into an average performance (e.g., average estimated accuracy as the mean fraction of correct predictions across all benchmark data sets). In both cases, method failure complicates the assessment and especially the comparison of performance of the methods under investigation such that the researcher finds themself wondering how to proceed adequately.
@@
The rest of this paper is structured as follows. In Section 2, we give an overview of the status quo by reviewing and summarizing the different manifestations and handlings of method failure reported in published comparison studies, followed by a detailed discussion of why discarding data or imputing values are usually inadequate when encountering method failure. Building on this, Section 3 is a compilation of alternative recommendations for handling method failure, which differentiates between different scopes of the study and research stages in which it may take place. In particular, we promote using fallback strategies when encountering method failure, which enables the aggregation of performance across data sets even when method failure occurs and on top of that, reflects the behavior of real method users across many settings. Considering possible time constraints, we also discuss “slimmed-down” ways in which authors can implement part of our recommendations if unable to follow them completely. Finally, in Section 4, we demonstrate our recommendations using two fictive comparison studies that illustrate the complexity of the discussed issues.
@@ 2.1 Status quo: occurrence and handling of method failure
Remove method Modify implementation Change software Discard data for failing methods Discard data for all methods Impute Performance measure
@@
Herrmann et al., (2021) Benchmark study on the predictive performance of thirteen survival time prediction methods based on eighteen high-dimensional multi-omics data.     * $\times$ *     $\times$ $\times$
Bischl et al., (2013) Comparison study of eleven classification methods on 49 real benchmark and simulated data sets.     * * *     $\times$ $\times$
Nießl et al., (2022)b Meta-study to investigate the effect of different imputation procedures on the performance assessment and ranking of the methods. Considered alternative procedures include imputation by the worst possible value and by a value that weighs the method’s overall proportion of method failure against its performance in the remaining folds.     * * *     $\times$
@@
    van Smeden et al van Smeden et al., (2016) observe that separated data sets (i.e., data sets leading to non-convergence) are commonly omitted from analyses.
@@
    Nießl et alNießl et al., (2022) perform an example benchmark study based on the results of the study by Herrmann et alHerrmann et al., (2021) to illustrate how different ways of imputation can influence the results of a benchmark study.
@@ 2.1.1 Common manifestations of method failure
Examples  Zapf et al., (2024) encounter non-convergence for two of the three methods under investigation. One returns an error, while the other outputs the current, non-converged estimate in these instances (Zapf et al., , 2024). Masaoud and Stryhn, (2010) encounter instances of “non-convergence or non-sensible estimates” (Masaoud and Stryhn, , 2010) for some method-data set combinations. In the study by Crowther et al., (2014), the methods are based on numerical integration and encounter estimation difficulties (non-convergence) with “challenging data sets(Crowther et al., , 2014). Hornung et al., (2024) report instances of method failure when methods are “not applicable” (Hornung et al., , 2024) to certain differences between training and test sets regarding the patterns of block-wise missingness. Additionally, they report calculation issues for one method explicitly disallowing certain structures in the training data set, resulting in failure. Calculation issues are reported by Ruxton and Neuhäuser, (2013), where some odds ratio estimation methods fail due to division by 00 when the $2\times 2$ contingency table contains one or more zero-entries. Another case of calculation issues can be found in the simulation study conducted by Dunias et al., (2024). For some data sets, the statistical learning method LASSO does not select any predictor variable, leading to a constant predictor and preventing the assessment of performance regarding calibration slope.
@@
Runtime issues  Another form of method failure often reported in comparison studies is excessive runtime. Some studies set fixed time budgets, so exceeding them is considered method failure. Others without fixed budgets encounter method failure when experiments for single methods run for an excessive amount of time (e.g. days), forcing the researcher to abort. Unlike non-convergence and memory issues, method failure is actively declared by the researchers when runtime issues occur, as without these (naturally necessary) limitations, a method might eventually generate an output at some point. Resulting implications are explored in Section 3.1.6.
@@ 2.1.2 Common approaches to handling method failure
Based on the published comparison studies, the handling of method failure can be commonly divided into two general approaches. Note that a literature review by Pawel et al., (2024), yielding a structured investigation of the frequencies of handling approaches from published simulation studies, results in four main handlings, which are among those we have commonly observed.
@@
Summarizing performance despite method failure Instead of resolving method failure, some authors summarize the methods’ performances despite existing failures. A widespread approach is to discard data sets associated with failure. The first version of this approach, consisting of discarding data sets for the failing methods only, is followed by Zapf et al., (2024) for the method affected by errors due to non-convergence. Fernández-Delgado et al., (2014), comparing the methods’ performance based on two performance measures, discard the associated data sets for the methods affected by method failure in the comparison based on average accuracy. The second version of discarding data, namely discarding the corresponding data sets for all methods, is applied by Hornung et al., (2024). Particularly, they argue against discarding the data sets for the failing methods only by saying that it would not yield a fair comparison since predictive performance is generally associated with the induced pattern of block-wise missingness (Hornung et al., , 2024).
@@
An alternative way of imputing is employed by Bischl et al., (2013), who impute performance values based on the proportion of repetitions in which the affected method fails. Below $20\%$, they impute by sampling from a normal distribution estimated from the remaining performance values of that model. Above $20\%$, on the other hand, they impute by the worst possible value with the reasoning that the method’s behavior is “too unreliable for the current data set” (Bischl et al., , 2013). A similar imputation strategy is employed by Herrmann et al., (2021). While using the same imputation rule when a method fails on more than $20\%$ of data sets, they otherwise impute by the mean performance of the method in the remaining “successful” data sets. In a subsequent project, Nießl et al., (2022) consider alternative procedures, including imputation by the worst possible value and by a value that weights the method’s overall proportion of method failure against its performance in the remaining folds. They show that the approach to imputing the undefined performance values has a notable effect on the performance ranking between the methods.
@@ 2.2.2 Why imputation is inadequate
We view it as mistaken to regard method failure as a form of missing data and have been careful to refer to “undefined” rather than “missing” values. The missing data literature defines missing values as those that exist and were intended to be recorded but for one reason or another are “not observed” (Rubin, , 2004; Carpenter et al., , 2023). Method failure is far closer to what is called a truncating event in the estimands literature (Kurland et al., , 2009; Kahan et al., , 2024), which denotes a value that does not exist. That is, when a method fails, the result we wished to record simply does not exist for that data set. Method failure should, therefore, not be viewed as obscuring the method’s results on a given data set. Instead, failure is the result we get! For this reason, we should be cautious about applying missing data approaches and particularly imputation to method failure. Note that this also includes measures of “weighted performance,” which weight the performance of a method according to the proportion of data sets with failure for that method. Using the correspondence between weighting and imputation approaches, this implies that weighting implicitly assigns some performance value in instances of method failure. The consequence of our dissuasion from “missing data-inspired” approaches is that our work focuses intensively on alternative ways to deal with method failure.
@@ 3 Principles and recommendations for handling method failure
The following section presents guidelines for handling method failure that are more appropriate than discarding data or imputation. Since, as previously mentioned, method failure can only be directly observed after the execution stage, the order of our recommendations aligns with the natural process of investigating these failures post-execution. However, we also highlight key considerations that can be addressed during the planning stage to help reduce both the potential for and unpredictability of failure in the first place.
@@ 3.1.1 Consider method failure as the result of a complex interplay
Besides influencing runtime and memory consumption (Zimmermann, , 2020), data set characteristics contributing to method failure are highlighted by the observation that a method usually fails for specific data sets, not all of them. Certain data characteristics can complicate or completely prevent required calculations of a method within the maximum number of iterations (non-convergence). Recall the study of van Smeden et al., (2016) addressing perfect separation of the outcome variable by one or more predictor variables as a cause for non-convergence of logistic regression algorithms. Additionally, Fernández-Delgado et al., (2014), in their extensive comparison study of supervised classification methods, list “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al., , 2014) as data set characteristics associated with method failure.
@@
In addition to the interplay of method and data set characteristics, however, software can also affect the (non-)failure of a method such that results are obtained in one software but not another. Consider, for instance, that Herrmann et al., (2021) reported that one of the methods under investigation was completely unusable due to an error in R that occurred using Windows as an operating system, while it did not occur under Ubuntu. The dependency on hardware, on the other hand, becomes clear in how working memory, besides being directly associated with possible memory issues, can also considerably affect runtimes.

### v3

@@ 1 Introduction
In methodological research, comparison studies aim to compare the performance of methods to provide empirical evidence on their behavior and help data analysts choose appropriate methods for their setting. To be reliable, they require careful planning of the design, execution, and reporting of the results. These issues have gained increasing attention in recent years, both in the context of simulation studies based on artificially generated data (Morris et al.,, 2019; Pawel et al., 2024b, ; Siepe et al.,, 2024) and in the context of real data-based benchmark studies (Boulesteix et al.,, 2017; Friedrich and Friede,, 2024).
@@
In this work, we focus on a particular aspect of the design of comparison studies that affects both simulation and benchmark studies, namely “method failure.” Method failure occurs when a method under investigation fails to produce an output for a (real or simulated) data set considered in the comparison study. The issue of method failure in comparison studies has begun to attract attention in recent years (Morris et al.,, 2019; Pawel et al., 2024b, ; Siepe et al.,, 2024), often focusing on non-convergence. However, there is little guidance on how to handle method failure appropriately, whether it occurs as non-convergence or in another form. In particular, published reviews have shown that method failure goes unreported in the majority of cases. For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur” (Hinds et al.,, 2018). Even more notably, only $12$ of $85$ applicable research articles reviewed by Morris et al., (2019) report convergence as a performance measure. This contrasts with our perception—both through our own experiences and informal discussions with colleagues—that the problem of method failure is highly prevalent. In particular, signs of method failure are often traceable in the code provided for reproducibility, even when not reported in the manuscript. This is often clear through the use of explicit error-handling mechanisms (e.g., functions try() or tryCatch() in R or try-except constructs in Python).
@@
To illustrate the difficulty of handling method failure adequately, imagine a researcher conducting a comparison study in statistics or predictive modeling to compare the performance of three methods based on four simulation/benchmark data sets. In the context of classical statistics, they likely focus on comparing the methods’ performances (e.g., bias) in “estimating one or more population quantities” (Morris et al.,, 2019), sometimes called “estimand(s),” and base the analysis on data simulated using a certain data-generating mechanism (DGM). In predictive modeling, they are likely to compare the methods’ predictive performances (e.g., accuracy), that is, their ability to make predictions on a new set of observations. Here, the analysis is typically performed on real benchmark data sets. For simplicity, we refer to both settings as “simulation studies” and “benchmark studies,” respectively (though admitting that there may be intermediate situations, e.g., simulation studies investigating predictive performance on artificial data).
@@
Av. accuracy ? 0.84 ?
Regardless of the form in which a researcher encounters method failure, it leads to NA for the associated method-data set combination where otherwise an estimate (for simulation studies) or a performance value (for benchmark studies) is expected. Table 1 depicts a corresponding summary of the results of the hypothetical researcher’s study across all simulation repetitions/benchmark data sets and methods. In simulation studies, the NAs complicate the derivation of performance (e.g., bias as the mean deviation of the estimates from the true value across all repetitions). In benchmark studies, on the other hand, they hinder the aggregation of performance values into an average performance (e.g., average estimated accuracy as the mean fraction of correct predictions across all benchmark data sets). In both cases, method failure complicates the assessment and especially the comparison of performance of the methods under investigation such that the researcher finds themself wondering how to proceed adequately.
@@
The rest of this paper is structured as follows. In Section 2, we give an overview of the status quo by reviewing and summarizing the different manifestations and handlings of method failure reported in published comparison studies, followed by a detailed discussion of why discarding data or imputing values is usually inadequate when encountering method failure. Building on this, Section 3 is a compilation of alternative recommendations for handling method failure, which differentiates between different scopes of the study and research stages in which it may take place. In particular, we promote using fallback strategies when encountering method failure, which enables the aggregation of performance across data sets even when method failure occurs, and on top of that, reflects the behavior of real method users across many settings. Considering possible time constraints, we also discuss “slimmed-down” ways in which authors can implement part of our recommendations if unable to follow them completely. Finally, in Section 4, we demonstrate our recommendations using two fictive comparison studies that illustrate the complexity of the discussed issues.
@@ 2.1 Status quo: occurrence and handling of method failure
Remove method Modify implementation Change software Discard data for failing methods Discard data for all methods Impute Include performance measure
@@
Herrmann et al., (2021) Benchmark study on the predictive performance of thirteen survival time prediction methods based on eighteen high-dimensional multi-omics data.     * $\times$ *     $\times$ b $\times$ $\times$
Bischl et al., (2013) Comparison study of eleven classification methods on 49 real benchmark and simulated data sets.     * * *     $\times$
Nießl et al., (2022)c Meta-study to investigate the effect of different imputation procedures on the performance assessment and ranking of the methods. Considered alternative procedures include imputation by the worst possible value and by a value that weighs the method’s overall proportion of method failure against its performance in the remaining folds.     * * *     $\times$
@@
    van Smeden et al., (2016) observe in published comparison studies that separated data sets (i.e., data sets leading to non-convergence) are commonly omitted from analyses.
@@
    Herrmann et al., (2021) apply a moderate form of data removal by restricting the analysis to the subset of smaller benchmark data sets for one method
  • c
    Nießl et al., (2022) perform an example benchmark study based on the results of the study by Herrmann et al., (2021) to illustrate how different ways of imputation can influence the results of a benchmark study.
@@ 2.1.1 Common manifestations of method failure
Examples Zapf et al., (2024) encounter non-convergence for two of the three methods under investigation. One returns an error, while the other outputs the current, non-converged estimate in these instances (Zapf et al.,, 2024). Masaoud and Stryhn, (2010) encounter instances of “non-convergence or non-sensible estimates” for some method-data set combinations. In the study by Crowther et al., (2014) the methods are based on numerical integration and encounter estimation difficulties (non-convergence) with “challenging data sets” (Crowther et al.,, 2014). Hornung et al., (2024) report instances of method failure when methods are “not applicable” to certain differences between training and test sets regarding the patterns of block-wise missingness. Additionally, they report calculation issues for one method explicitly disallowing certain structures in the training data set, resulting in failure. Calculation issues are also reported in the study by Ruxton and Neuhäuser, (2013), where some odds ratio estimation methods fail due to division by 0 when the $2\times 2$ contingency table contains one or more zero entries. Fernández-Delgado et al., (2014) report several data set characteristics associated with issues in required calculations, including “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al.,, 2014). Additionally, a group of methods in their study requires a certain minimum number of observations per outcome class and is therefore not applicable to particularly small data sets. Another case of calculation issues can be found in the simulation study conducted by Dunias et al., (2024). For some data sets, the statistical learning method LASSO does not select any predictor variable, leading to a constant predictor and preventing the assessment of performance regarding calibration slope.
@@
Runtime issues Another form of method failure often reported in comparison studies is excessive runtime. Some studies set fixed time budgets, so exceeding them is considered method failure. Others without fixed budgets encounter method failure when experiments for single methods run for an excessive amount of time (e.g., days), forcing the researcher to abort. Unlike non-convergence and memory issues, method failure is actively declared by the researchers when runtime issues occur, as without these (naturally necessary) limitations, a method might eventually generate an output at some point. Resulting implications are discussed in Section 3.1.6.
@@ 2.1.2 Common approaches to handling method failure
Based on the published comparison studies, the handling of method failure can be commonly divided into two general approaches. Note that a literature review by Pawel et al., 2024a , yielding a structured investigation of the frequencies of handling approaches from published simulation studies, results in four main handlings, which are among those we have commonly observed.
@@
Summarizing performance despite method failure Instead of resolving method failure, some authors summarize the methods’ performances despite existing failures. A widespread approach is to discard data sets associated with failure. The first version of this approach, consisting of discarding data sets for the failing methods only, is followed by Zapf et al., (2024) for the method affected by errors due to non-convergence. Fernández-Delgado et al., (2014), comparing the methods’ performances based on two performance measures, discard the associated data sets for the methods affected by method failure in the comparison based on average accuracy. The second version of discarding data, namely discarding the corresponding data sets for all methods, is applied by Hornung et al., (2024). Particularly, they argue against discarding the data sets for the failing methods only by saying that it would not yield a fair comparison since predictive performance is generally associated with the induced pattern of block-wise missingness (Hornung et al.,, 2024).
@@
An alternative way of imputing is employed by Bischl et al., (2013), who impute performance values based on the proportion of repetitions in which the affected method fails. Below $20\%$, they impute by sampling from a normal distribution estimated from the remaining performance values of that model. Above $20\%$, on the other hand, they impute by the worst possible value with the reasoning that the method’s behavior is “too unreliable for the current data set” (Bischl et al.,, 2013). A similar imputation strategy is employed Herrmann et al., (2021). While using the same imputation rule when a method fails on more than $20\%$ of data sets, they otherwise impute by the mean performance of the method in the remaining “successful” data sets. In a subsequent project, Nießl et al., (2022) consider alternative procedures, including imputation by the worst possible value and by a value that weights the method’s overall proportion of method failure against its performance in the remaining folds. They show that the approach to imputing the undefined performance values has a notable effect on the performance ranking between the methods.
@@ 2.2.2 Why imputation is inadequate
We view it as mistaken to regard method failure as a form of missing data and have been careful to refer to “undefined” rather than “missing” values. The missing data literature defines missing values as those that exist and were intended to be recorded but for one reason or another are “not observed” (Rubin,, 2004; Carpenter et al.,, 2023). Method failure is far closer to what is called a truncating event in the estimands literature (Kurland et al.,, 2009; Kahan et al.,, 2024), which denotes a value that does not exist. That is, when a method fails, the result we wished to record simply does not exist for that data set. Method failure should, therefore, not be viewed as obscuring the method’s results on a given data set. Instead, failure is the result we get! For this reason, we should be cautious about applying missing data approaches, and particularly imputation, to method failure. Note that this also includes measures of “weighted performance,” which weight the performance of a method according to the proportion of data sets with failure for that method. Using the correspondence between weighting and imputation approaches, this implies that weighting implicitly assigns some performance value in instances of method failure. The consequence of our dissuasion from “missing data-inspired” approaches is that our work focuses intensively on alternative ways to deal with method failure.
@@ 3 Principles and recommendations for handling method failure
The following section presents guidelines for handling method failure that are more appropriate than discarding data or imputation. Since, as previously mentioned, method failure can only be directly observed after the execution stage, the order of our recommendations aligns with the natural process of investigating these failures post-execution. However, we also highlight key considerations that can be addressed during the planning stage to help reduce both the potential for and unpredictability of failure in the first place. An overview of the recommendations can be found in Table 4.
@@ 3.1 Practical recommendations for the handling of method failure
Table 4: Checklist of recommendations for handling method failure.
Recommended error handling workflow Section
1. Investigate and understand the interplay underlying each instance of method failure
Interplay may consist of data set characteristics, the method’s implementation, the given hardware, and software.
3.1.1
2. Check for the correct use of the method
Distinguish between “use beyond original scope” and “method misuse” when detecting that a failing method is used differently than intended by its developers.
3.1.2
3. Check if data matches the scope of the study
Check if the characteristics of the data associated with failure (and the failure itself) are realistic for the targeted users.
3.1.3
4. Be realistic when modifying method implementations
For early-stage research. Restrict yourself to parameter adjustments that are realistic for future targeted users.
3.1.4
5. Consider a fallback strategy
For late-stage research. Evaluate the performance of method pipelines of the type “use method A when it successfully produces an output, use method B otherwise.” Define the pipelines based on preliminary considerations and avoid selectively reporting the “best” pipelines.
3.1.5
6. Special case: runtime issues
When real users running the affected method only on a single data set can be assumed to successfully obtain a model output, run the method on a preselected random subset of all data sets (selection best made in the planning stage).
3.1.6
7. Report failure proportions, underlying interplays, applied approaches, and share code
Provide valuable insights to real users into the methods’ usability and enhance a deeper understanding of method failure in general.
3.1.7
8. Think ahead and pre-register, but remain flexible
Conduct a pilot study before pre-registering the handling of method failure. Be aware that unforeseen failures may still occur; therefore, include error-handling mechanisms in your code from the start.
3.1.8
Slimmed-down error handling workflow when time constraints prevent recommendations 1-8
Report (i) performance results after discarding the data sets for all methods, (ii) performance results after discarding the data sets for failing methods only, and (iii) failure proportions for all methods. Note that this approach does not provide a comparison of unconditional performance. 3.2
@@ 3.1.1 Consider method failure as the result of a complex interplay
Besides influencing runtime and memory consumption (Zimmermann,, 2020), data set characteristics contributing to method failure are highlighted by the observation that a method usually fails for specific data sets, not all of them. Certain data characteristics can complicate or completely prevent required calculations of a method within the maximum number of iterations (non-convergence), time, or memory budgets. Note that this association is addressed in several comparison studies listed in Section 2.1.1. For instance, the study by van Smeden et al., (2016) addresses perfect separation of the outcome variable by one or more predictor variables as a cause of non-convergence of logistic regression algorithms. Recall also the data set characteristics associated with failure that are reported by Fernández-Delgado et al., (2014), including“collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al.,, 2014).
@@
In addition to the interplay of method and data set characteristics, however, software can also affect the (non-)failure of a method such that results are obtained in one software but not another. Consider, for instance, that Herrmann et al., (2021) report that one of the methods under investigation is completely unusable due to an error in R that occurred using Windows as an operating system, while it does not occur under Ubuntu. The dependency on hardware, on the other hand, becomes clear in how working memory, besides being directly associated with possible memory issues, can also considerably affect runtimes.